# Analysis and Synthesis Denoisers

### v2

@@ 2.2 FB-PnP algorithm with approximated AD
find $\text{find }\widehat{x}_{\mathcal{A}}\in\underset{x\in\mathbb{R}^{N}}{\operatorname{argmin}}\;\frac{1}{2}\|Ax-y\|^{2}+g_{\lambda}(\Gamma x).$ (2.6)
@@
$\begin{array}[]{l}x_{0}\in\mathbb{R}^{N},z_{0}\in\mathbb{R}^{S},\\ \text{for }k=0,1,\ldots\\ \left\lfloor\begin{array}[]{l}{v}_{k}=x_{k}-\tau A^{*}(Ax_{k}-y),\\ u_{k+1}={G}^{\mathcal{A}}_{L,\tau\lambda,v_{k}}(u_{k}),\\ x_{k+1}=v_{k}-\Gamma u_{k+1},\end{array}\right.\end{array}$ (2.7)
@@
The following convergence result directly follows from Theorem 2.1 and [17, Thm. 3.4].
@@
Let ${(x_{k})}_{k\in\mathbb{N}}$ be generated by algorithm (2.7). Assume that there exist $\underline{\tau}>0$ and $\underline{\sigma}>0$ such that $\underline{\tau}\leq\tau\leq(2-\underline{\tau})\|A\|_{S}^{-2}$ and $\underline{\sigma}\leq\sigma\leq(2-\underline{\sigma})\|\Gamma\|_{S}^{-2}$. When $L\to\infty$, i.e. when we use an infinite number of iterations in Model 2.2, then ${(x_{k})}_{k\in\mathbb{N}}$ converges to a solution to problem (2.6).
Theorem 2.4 only holds when $L\to\infty$, i.e. when $G^{\mathcal{A}}$ has a very large number of sub-iterations. In practice, this is often intractable and only a fixed finite number of iterations are used. For instance, in the context of DDL, typically $L$ will be smaller than $20$. We propose to investigate to what extent this impacts the convergence and the nature of the final solution. Note that in algorithm (2.7), the unrolled denoiser benefits from a warm restart using the output $u_{k}$ of the inner iterations $\widetilde{G}^{\mathcal{A}}$ obtained at iteration $k$. We will take advantage of this warm restart to show the convergence of ${(x_{k})}_{k\in\mathbb{N}}$ when $L=1$. In this particular case, at iteration $k\in G$, the update of $u_{k}$ in algorithm (2.7) boils down to
$u_{k+1}={G}^{\mathcal{A}}_{1,\tau\lambda,v_{k}}(u_{k})=T^{\mathcal{A}}_{\tau\lambda,v_{k}}(u_{k}),$ (2.8)
@@
In the following result, we show that algorithm (2.7) with $L=1$ corresponds to the scaled primal-dual algorithm proposed by [35] (see Section 6 in this article, equation (40)). We can then deduce convergence guarantees for ${(x_{k})}_{k\in\mathbb{N}}$, leveraging [35, Thm. 1].
@@
Let ${(x_{k})}_{k\in\mathbb{N}}$ and ${(u_{k})}_{k\in\mathbb{N}}$ be sequences generated by algorithm (2.7) with $L=1$. Assume that there exist $\underline{\tau}>0$ and $\underline{\sigma}>0$ such that $\underline{\tau}\leq\tau\leq(2-\underline{\tau})\|A\|_{S}^{-2}$ and $\underline{\sigma}\leq\sigma\leq(2-\underline{\sigma})\|\Gamma\|_{S}^{-2}$. Then the following statements hold:
@@
    ${(x_{k},\tau^{-1}u_{k})}_{k\in\mathbb{N}}$ converges to a solution to the saddle-point problem
    $\underset{x\in\mathbb{R}^{N}}{\operatorname{minimize}}\;\max_{u\in\mathbb{R}^{S}}\;\frac{1}{2}\|Ax-y\|^{2}+\left\langle\Gamma x,u\right\rangle-g_{\lambda}^{*}(u).$ (2.9)
@@
$u_{k+1}$ $=\operatorname{prox}_{\sigma g_{\lambda\tau}^{*}}\Big{(}u_{k}-\sigma\Gamma^{*}\Gamma u_{k}+\sigma\Gamma^{*}v_{k}\Big{)}$
$=\operatorname{prox}_{\sigma g_{\lambda\tau}^{*}}\Big{(}u_{k}-\sigma\Gamma^{*}\Gamma u_{k}+\sigma\Gamma^{*}\big{(}x_{k}-\tau A^{*}(Ax_{k}-y)\big{)}\Big{)}$
$=\operatorname{prox}_{\sigma g_{\lambda\tau}^{*}}\Big{(}u_{k}+\sigma\Gamma^{*}\Big{(}x_{k}-\tau A^{*}\big{(}Ax_{k}-y\big{)}-\Gamma u_{k}\Big{)}\Big{)}.$
@@
$\begin{array}[]{l}\text{for }k=0,1,\ldots\\ \left\lfloor\begin{array}[]{l}u_{k+1}=\operatorname{prox}_{\sigma g_{\lambda\tau}^{*}}\Big{(}u_{k}+\sigma\Gamma^{*}\Big{(}x_{k}-\tau A^{*}\big{(}Ax_{k}-y\big{)}-\Gamma u_{k}\Big{)}\Big{)}\\ x_{k+1}=x_{k}-\tau A^{*}(Ax_{k}-y)-\Gamma u_{k+1}.\end{array}\right.\end{array}$ (2.10)
Further, noticing that $g_{\lambda\tau}=\tau g_{\lambda}$, and applying subsequently Prop. 24.8(v) and Prop. 13.23(i) from [4], we have
$(\forall u\in\mathbb{R}^{S})\quad\operatorname{prox}_{\sigma g_{\lambda\tau}^{*}}(u)=\operatorname{prox}_{\sigma\tau g_{\lambda}^{*}}(u)=\operatorname{prox}_{\sigma\tau g_{\lambda}^{*}(\tau^{-1}\cdot)}(u)=\tau\operatorname{prox}_{\sigma\tau^{-1}g_{\lambda}^{*}}(\tau^{-1}u).$
@@
$\begin{array}[]{l}\text{for }k=0,1,\ldots\\ \left\lfloor\begin{array}[]{l}u_{k+1}=\tau\operatorname{prox}_{\sigma\tau^{-1}g_{\lambda}^{*}}\bigg{(}\tau^{-1}\Big{(}u_{k}+\sigma\Gamma^{*}\Big{(}x_{k}-\tau A^{*}\big{(}Ax_{k}-y\big{)}-\Gamma u_{k}\Big{)}\Big{)}\bigg{)}\\ x_{k+1}=x_{k}-\tau A^{*}(Ax_{k}-y)-\Gamma u_{k+1}.\end{array}\right.\end{array}$ (2.11)
@@
$\begin{array}[]{l}\text{for }k=0,1,\ldots\\ \left\lfloor\begin{array}[]{l}\widetilde{u}_{k+1}=\operatorname{prox}_{\frac{\sigma}{\tau}g_{\lambda}^{*}}\Big{(}\widetilde{u}_{k}+\frac{\sigma}{\tau}\Gamma^{*}\big{(}x_{k}-\tau A^{*}(Ax_{k}-y)-\tau\Gamma\widetilde{u}_{k}\big{)}\Big{)}\\ x_{k+1}=x_{k}-\tau A^{*}(Ax_{k}-y)-\tau\Gamma\widetilde{u}_{k+1}.\end{array}\right.\end{array}$ (2.12)
By noticing that algorithm (2.12) corresponds to the scaled primal-dual algorithm proposed by [35] (see Section 6 in this article, equation (40)), [35, Thm. 1] can directly be applied to deduce the convergence results, using the scaled version of the algorithm given in Section 6 of the same article.
@@
In the particular case $g=\lambda\|\cdot\|_{1}$, then Theorem 2.5 shows that
@@
where $\mathcal{F}_{\lambda}^{\mathcal{A}}(\Gamma)=\big{\{}x\ |\ \inf\{\|z\|_{\infty}\ |\ \Gamma z=x\}\leq\lambda\big{\}}$.
@@ 3.1 Synthesis problem formulations
find $\text{find }\quad u^{\dagger}_{D}=\operatorname*{arg\,min}_{u\in\mathbb{R}^{N}}\frac{1}{2}\|u-v\|^{2}+g_{\lambda}^{*}(D^{*}u)=\operatorname{prox}_{g_{\lambda}^{*}\circ D^{*}}(v).$ (3.4)
@@
$-u^{\dagger}_{D}\in\partial_{\frac{1}{2}\|\cdot-v\|^{2}_{2}}(Dz^{\dagger}_{D})=\{Dz^{\dagger}_{D}-v\}\quad\Leftrightarrow\quad Dz^{\dagger}_{D}=v-u^{\dagger}_{D}.$ (3.5)
@@
Is it important to note that here, we don’t consider the traditional output of the synthesis formulation $z_{S}^{\dagger}$ but the denoised image $x_{S}^{\dagger}$ that is recovered from it. While $z_{S}^{\dagger}$ might be non-unique and the application linking some input $v$ to $z_{S}^{\dagger}(v)=\operatorname{prox}_{g_{\lambda}}(v)$ is non-smooth, it is not the case for $x_{S}^{\dagger}$, which results from the proximity operator. We can thus deduce from Proposition 3.1 that the PnP-FB algorithm (1.5) with $G$ being the SD aims to
find $\text{find }\widehat{x}_{\mathcal{S}}\in\underset{x\in\mathbb{R}^{N}}{\operatorname{argmin}}\;\frac{1}{2}\|Ax-y\|^{2}+D\rhd g_{\lambda}(x).$ (3.6)
@@
find $\text{find }\widehat{z}_{\mathcal{S}}\in\underset{z\in\mathbb{R}^{S}}{\operatorname{Argmin}}\;\frac{1}{2}\|ADz-y\|^{2}+g_{\lambda}(z).$ (3.7)
@@
$\begin{array}[]{l}\text{for }k=0,1,\ldots\\ \left\lfloor\begin{array}[]{l}v_{k}=z_{t}-\gamma D^{*}A^{*}(ADz_{t}-y)),\\ z_{k+1}=\operatorname{prox}_{\gamma g_{\lambda}}(v_{k}),\end{array}\right.\end{array}$ (3.8)
where $z_{0}\in\mathbb{R}^{S}$, and $\gamma>0$ is the step-size chosen to enable convergence of ${(z_{k})}_{k\in\mathbb{N}}$ to a solution to (3.7[17, Thm. 3.4].
The result below highlights the links between the full synthesis problem (3.6) and the analysis problem with SD (3.7).
@@
where in (3.10), the Fenchel-Rockafellar strong duality is applied to $f^{*}$ and $(g_{\lambda}^{*}\circ D^{*})$, with $f^{**}=f$ as $f\in\Gamma_{0}(\mathbb{R}^{N})$, and the last equality is obtained using (1.9). This shows (3.9), i.e., that the values of the two problems minima are the same.
@@
Let $\widehat{x}_{\mathcal{S}}$ be a solution to (3.6). Combining (3.12) and (3.13), we obtain
$f(D\widehat{z}_{\mathcal{S}})+D\rhd g_{\lambda}(D\widehat{z}_{\mathcal{S}})\leq f(\widehat{x}_{\mathcal{S}})+D\rhd g_{\lambda}(\widehat{x}_{\mathcal{S}}).$
Hence $\widehat{x}_{\mathcal{S}}=D\widehat{z}_{\mathcal{S}}$.
(ii) Let $\widehat{x}_{\mathcal{S}}$ be a solution to (3.6). We will show that there exist some $\widehat{z}_{\mathcal{S}}$ such that $\widehat{x}_{\mathcal{S}}=D\widehat{z}_{\mathcal{S}}$ and $\widehat{z}_{\mathcal{S}}$ solution to (3.7). Assume that the component of $\widehat{x}_{\mathcal{S}}$ contained in ${\mathrm{im}(D)}^{\bot}=\ker D^{*}$ is not 00, then
$\sup_{v\in\mathbb{R}^{N}}\langle v,\widehat{x}_{\mathcal{S}}\rangle-g_{\lambda}^{*}(D^{*}v)\geq\sup_{v\in\ker D^{*}}\langle v,\widehat{x}_{\mathcal{S}}\rangle-g_{\lambda}^{*}(0)\geq+\infty.$
As the value of (3.10) is finite, this would violate the optimality of $\widehat{x}_{\mathcal{S}}$. Thus $\widehat{x}_{\mathcal{S}}$ is contained in the image of $D$, and there indeed exists $\widehat{z}_{\mathcal{S}}\in\mathbb{R}^{S}$ such that $\widehat{x}_{\mathcal{S}}=D\widehat{z}_{\mathcal{S}}$. By definition of the infimal post-composition, we have
$f(\widehat{x}_{\mathcal{S}})+D\rhd g_{\lambda}(\widehat{x}_{\mathcal{S}})$ $=f(\widehat{x}_{\mathcal{S}})+\inf_{z\in\mathbb{R}^{S},\widehat{x}_{\mathcal{S}}=Dz}g_{\lambda}(z)$
$=\min_{z\in\mathbb{R}^{S},\widehat{x}_{\mathcal{S}}=Dz}f(Dz)+g_{\lambda}(z).$ (3.14)
@@
Hence, thanks to (3.11), $\widehat{z}_{\mathcal{S}}$ is a minimizer of (3.7) and this conclude the proof.
@@
In the particular case when $g=\lambda\|\cdot\|_{1}$, then Theorem 3.2 ensures that

### v3

@@ 2.2 FB-PnP algorithm with approximated AD
$\text{find }\widehat{x}_{\mathcal{A}}\in\underset{x\in\mathbb{R}^{N}}{\operatorname{Argmin}}\;\frac{1}{2}\|Ax-y\|^{2}+g_{\lambda}(\Gamma x).$ (2.6)
@@
[…]
@@
The following convergence result directly follows from Theorem 2.1 and [20, Thm. 3.4].
@@
Let ${(x_{k})}_{k\in\mathbb{N}}$ be generated by algorithm (2.7). Assume that $0<\tau<2\|A\|_{S}^{-2}$ and $0<\sigma<2\|\Gamma\|_{S}^{-2}$. When $L\to\infty$ in Model 2.2 (i.e., when the proximity operator in (1.6) is computed accurately) then ${(x_{k})}_{k\in\mathbb{N}}$ converges to a solution to problem (2.6).
Theorem 2.4 is purely theoretical as it only holds when $L\to\infty$, i.e. when $\widetilde{G}^{\mathcal{A}}$ has a very large number of sub-iterations. In practice, this is intractable and only a fixed finite number of iterations are used. For instance, in the context of DDL, typically $L$ will be smaller than $20$. We propose to investigate to what extent this impacts the convergence and the nature of the final solution. Note that in algorithm (2.7), the unrolled denoiser benefits from a warm restart using the output $u_{k}$ of the inner iterations $\widetilde{G}^{\mathcal{A}}$ obtained at iteration $k$. We will take advantage of this warm restart to show the convergence of ${(x_{k})}_{k\in\mathbb{N}}$ when $L=1$. In this particular case, at iteration $k\in\mathbb{N}$, the update of $u_{k}$ in algorithm (2.7) boils down to
$u_{k+1}=\widetilde{G}^{\mathcal{A}}_{1,\tau\lambda,v_{k}}(u_{k})=T^{\mathcal{A}}_{\tau\lambda,v_{k}}(u_{k}),$ (2.8)
@@
In the following result, we show that algorithm (2.7) with $L=1$ corresponds to the scaled primal-dual algorithm proposed by [38] (see Section 6 in this article, equation (40)). We can then deduce convergence guarantees for ${(x_{k})}_{k\in\mathbb{N}}$, leveraging [38, Thm. 1].
@@
Let ${(x_{k})}_{k\in\mathbb{N}}$ and ${(u_{k})}_{k\in\mathbb{N}}$ be sequences generated by algorithm (2.7) with $L=1$. Assume that $0<\tau<2\|A\|_{S}^{-2}$ and $0<\sigma<2\|\Gamma\|_{S}^{-2}$. Then the following statements hold:
@@
    ${(x_{k},\tau^{-1}u_{k})}_{k\in\mathbb{N}}$ converges to a solution to the saddle-point problem aiming to
    $\text{find }(\widehat{x}_{\mathcal{A}},\widehat{u}_{\mathcal{A}})\in{\operatorname{Arg}}\min_{x\in\mathbb{R}^{N}}\max_{u\in\mathbb{R}^{S}}\;\frac{1}{2}\|Ax-y\|^{2}+\left\langle\Gamma x,u\right\rangle-g_{\lambda}^{*}(u).$ (2.9)
@@
$u_{k+1}$ $=\operatorname{prox}_{\sigma g_{\lambda\tau}^{*}}\Big{(}u_{k}-\sigma\Gamma\Gamma^{*}u_{k}+\sigma\Gamma v_{k}\Big{)}$
$=\operatorname{prox}_{\sigma g_{\lambda\tau}^{*}}\Big{(}u_{k}-\sigma\Gamma\Gamma^{*}u_{k}+\sigma\Gamma\big{(}x_{k}-\tau A^{*}(Ax_{k}-y)\big{)}\Big{)}$
$=\operatorname{prox}_{\sigma g_{\lambda\tau}^{*}}\Big{(}u_{k}+\sigma\Gamma\Big{(}x_{k}-\tau A^{*}\big{(}Ax_{k}-y\big{)}-\Gamma^{*}u_{k}\Big{)}\Big{)}.$
@@
[…]
Further, noticing that $g_{\lambda\tau}=\tau g_{\lambda}$ (see definition of $g_{\lambda}$ after (1.6)), and applying subsequently Prop. 24.8(v) and Prop. 13.23(i) from [4], we have
$(\forall u\in\mathbb{R}^{S})\quad\operatorname{prox}_{\sigma g_{\lambda\tau}^{*}}(u)=\operatorname{prox}_{\sigma\tau g_{\lambda}^{*}}(u)=\tau\operatorname{prox}_{\sigma\tau^{-1}g_{\lambda}^{*}}(\tau^{-1}u).$
@@
[…]
@@
[…]
which is equivalent to the scaled Loris-Verhoeven primal-dual algorithm proposed by [38] (see Appendix A). Hence [38, Thm. 1] can directly be applied to deduce the convergence results.
@@
In the particular case $g_{\lambda}=\lambda\|\cdot\|_{1}$, then Theorem 2.5 shows that
@@
where $\mathcal{F}_{\lambda}^{\mathcal{A}}(\Gamma)=\big{\{}x\ |\ \inf\{\|z\|_{\infty}\ |\ \Gamma^{*}z=x\}\leq\lambda\big{\}}$. This is true for any choice of $\Gamma$, including wavelet transforms or the finite difference operator (so that $g_{\lambda}\circ\Gamma$ corresponds to the anisotropic total variation).
@@ 3 Synthesis denoiser for FB-PnP algorithm
In this section, we use the subscript $D$ in $x^{\dagger}_{D}$ and $z^{\dagger}_{D}$ to refer to outputs of denoising tasks involving a fixed synthesis dictionary $D$, from which the SD is built. Similarly to Section 2, we further use the subscript $\mathcal{S}$ to denote solutions associated with the synthesis formulation of the full inverse problem, where $D$ is used to reconstruct the signal from sparse codes. This notational distinction aims to separate the denoising phase from the inverse problem phase.
@@ 3.1 Synthesis problem formulations
The following result highlights that the synthesis denoising problem (3.1) corresponds to a proximity operator. This fact has also been noted in [12, Sect. 5.3] and [8, Prop. 1] to reformulate the ADMM and the Douglas-Rachford iterations, respectively.
@@
$\text{find }\quad u^{\dagger}_{D}=\underset{u\in\mathbb{R}^{N}}{\operatorname{argmin}}\;\frac{1}{2}\|u-v\|^{2}+g_{\lambda}^{*}(D^{*}u)=\operatorname{prox}_{g_{\lambda}^{*}\circ D^{*}}(v).$ (3.4)
@@
$u^{\dagger}_{D}\in\partial_{\frac{1}{2}\|\cdot-v\|^{2}_{2}}(Dz^{\dagger}_{D})=\{Dz^{\dagger}_{D}-v\}\quad\Leftrightarrow\quad Dz^{\dagger}_{D}=v-u^{\dagger}_{D}.$ (3.5)
@@
Is it important to note that here, we don’t consider the traditional output of the synthesis formulation $z^{\dagger}_{D}$ but the denoised image $x^{\dagger}_{D}$ that is recovered from it. While $z^{\dagger}_{D}$ might be non-unique and the application linking some input $v$ to $z^{\dagger}_{D}$ is non-smooth, it is not the case for $x^{\dagger}_{D}$, which results from the proximity operator. We can thus deduce from Proposition 3.1 that the PnP-FB algorithm (1.5) with $G$ being the SD aims to
$\text{find }\widehat{x}_{\mathcal{S}}\in\underset{x\in\mathbb{R}^{N}}{\operatorname{Argmin}}\;\frac{1}{2}\|Ax-y\|^{2}+D\rhd g_{\lambda}(x).$ (3.6)
@@
$\text{find }\widehat{z}_{\mathcal{S}}\in\underset{z\in\mathbb{R}^{S}}{\operatorname{Argmin}}\;\frac{1}{2}\|ADz-y\|^{2}+g_{\lambda}(z).$ (3.7)
@@
[…]
where $z_{0}\in\mathbb{R}^{S}$, and $\gamma>0$ is the step-size chosen to enable convergence of ${(z_{k})}_{k\in\mathbb{N}}$ to a solution to (3.7[20, Thm. 3.4].
The result below highlights the links between the full synthesis problem (3.7) and the analysis problem with SD (3.6). Similar results can be found in [8, Prop. 1] to deduce an alternative formulation for the Douglas-Rachford algorithm.
@@
where in (3.10), the Fenchel-Rockafellar strong duality is applied to $f^{*}$ and $(g_{\lambda}^{*}\circ D^{*})$, with $f^{**}=f$ as $f\in\Gamma_{0}(\mathbb{R}^{N})$, and the last equality is obtained using [4, Prop. 13.24]. This shows (3.9), i.e., that the values of the two problems minima are the same.
@@
Combining (3.12) and (3.13), we obtain $f(D\widehat{z}_{\mathcal{S}})+D\rhd g_{\lambda}(D\widehat{z}_{\mathcal{S}})\leq\min_{x\in\mathbb{R}^{N}}f(x)+D\rhd g_{\lambda}(x).$
Hence $D\widehat{z}_{\mathcal{S}}\in\underset{x\in\mathbb{R}^{N}}{\operatorname{Argmin}}\;f(x)+D\rhd g_{\lambda}(x)$, so $\widehat{x}_{\mathcal{S}}=D\widehat{z}_{\mathcal{S}}$ is solution to (3.6).
(ii) Let $\widehat{x}_{\mathcal{S}}$ be a solution to (3.6). We will show that there exist some $\widehat{z}_{\mathcal{S}}$ such that $\widehat{x}_{\mathcal{S}}=D\widehat{z}_{\mathcal{S}}$ and $\widehat{z}_{\mathcal{S}}$ solution to (3.7). Assume that the component of $\widehat{x}_{\mathcal{S}}$ contained in ${\mathrm{im}(D)}^{\bot}=\ker D^{*}$ is not 0, then $\sup_{v\in\mathbb{R}^{N}}\langle v,\widehat{x}_{\mathcal{S}}\rangle-g_{\lambda}^{*}(D^{*}v)\geq\sup_{v\in\ker D^{*}}\langle v,\widehat{x}_{\mathcal{S}}\rangle-g_{\lambda}^{*}(0)\geq+\infty$. As the value of (3.10) is finite, this would violate the optimality of $\widehat{x}_{\mathcal{S}}$. Thus $\widehat{x}_{\mathcal{S}}$ is contained in the image of $D$, and there indeed exists $\widehat{z}_{\mathcal{S}}\in\mathbb{R}^{S}$ such that $\widehat{x}_{\mathcal{S}}=D\widehat{z}_{\mathcal{S}}$. By definition of the infimal post-composition, we have
$f(\widehat{x}_{\mathcal{S}})+D\rhd g_{\lambda}(\widehat{x}_{\mathcal{S}})=f(\widehat{x}_{\mathcal{S}})+\inf_{z\in\mathbb{R}^{S},\widehat{x}_{\mathcal{S}}=Dz}g_{\lambda}(z)=\min_{z\in\mathbb{R}^{S},\widehat{x}_{\mathcal{S}}=Dz}f(Dz)+g_{\lambda}(z).$ (3.14)
@@
Hence, thanks to (3.11), $\widehat{z}_{\mathcal{S}}$ is a minimizer of (3.7) and this concludes the proof.
@@
In the particular case when $g_{\lambda}=\lambda\|\cdot\|_{1}$, then Theorem 3.2 ensures that